OpenStudy (anonymous):

A 10.0 Ω resistor is in parallel with a 30.0 Ω resistor and a 5.0 Ω resistor. What is the equivalent resistance of this circuit? A. 0.02 Ω B. 0.33 Ω C. 3.0 Ω D. 45 Ω

OpenStudy (michele_laino):

please, wait a moment

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

the requested resistance is: \[\Large R = \frac{1}{{\frac{1}{{10}} + \frac{1}{{30}} + \frac{1}{5}}}\]

OpenStudy (anonymous):

we get 3? so our solution is choice C ?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay!! thanks!

OpenStudy (michele_laino):

:)